Diffusion Morning Report 04/14/11

Ted Berg tberg at stanford.edu
Thu Apr 14 06:27:52 PDT 2011


Good Morning all,
     Looking good this morning. PH3 for Epi is empty. Trying to 
expedite. Tylan Nitride and Thermco Nitride need tests after DCS 
restored. Monitoring Tylan LTO O2 flow. Thermco 2 no TLC flow.  4160m PM 
completed.No other issues at this time.

					
	Date: 	4/14/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*5.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*1.94/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*.34/30* 	** 	**
Tylan 10-Nitride 	*yellow* 	*needs test after DCS restored* 
*11.03/30* 	** 	*2*
Tylan 11-LTO 	*yellow* 	*O2 flow issues* 	*9.67/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*down* 	*needs test after DCS restored* 	** 	** 	*1*
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*down* 	*no PH3* 	** 	** 	**
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 449 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110414/2b26d254/attachment.html>


More information about the diffusion mailing list